# CodePMP: Scalable Preference Model Pretraining for Large Language Model Reasoning

### v1

@@ Appendix A Hyperparameters
We outline key hyperparameters in this Table 3. In the tables, WSD refers to the warmup-stable-decay learning rate scheduler( (Hu et al., 2024)), which has the benefit of reducing the time required for scaling law experiments.
Table 3: Hyperparameters for CodePMP training, mathematical reasoning RM finetuning, and logical reasoning RM finetuning.
CodePMP Mathematical RM Logical RM
HP 1.5B 7B 1.5B 7B 1.5B 7B
epoch 1 1 1 1 1 1
bs 1024 1024 64 64 64 64
lr 3e-6 1e-6 1e-6 3e-7 1e-5 1e-5
lr scheduler WSD WSD WCD WCD WCD WCD
warmup ratio 0.03 0.03 0.03 0.03 0.25 0.25
decay ratio 0.1 0.1 - - - -
weight decay 0.1 0.1 0 0 0 0
max length 1024 1024 1024 1024 1024 1024
Appendix B Training Pipeline
Figure B presents an overview of the complete training pipeline. The process begins with base language model (LM) pretraining on trillions of tokens from general text, followed by a preference model pretraining (PMP) phase using billions of tokens from code preference pairs. Finally, the model is finetuned on a smaller, more specialized dataset relevant to reasoning tasks, typically consisting of millions of tokens.
Figure 9: An overview of the complete training pipeline.
Appendix C CodePMP Dataset
As shown in Table 4, the constructed CodePMP dataset consists of a total of 28 million files, accounting for 19 billion tokens across various languages. The dataset is primarily composed of Python files, with 20 million files and 13.1 billion tokens, followed by Notebook files, contributing 3 million files and 2.1 billion tokens, and other programming languages with 5 million files and 3.8 billion tokens. This diverse dataset supports the pretraining phase, aiding the model in generalizing across multiple reasoning tasks.
The average length of chosen responses varies slightly between different languages. Python files exhibit an average response length of 170 tokens for chosen responses and 167 tokens for rejected responses, while Notebook files have slightly shorter average lengths of 158 and 155.5 tokens, respectively. Other languages show the highest average response lengths, with 213.2 tokens for chosen and 210 tokens for rejected responses. Overall, the average length of chosen responses across the dataset is 194.5 tokens, while the average length of rejected responses is 189.9 tokens, indicating minimal bias in the dataset based on response length.
Table 4: Amount and average length statistics of CodePMP dataset.
Amount Statistics Average Length
Language Files (M) Tokens (B) Chosen Rejected
Python 20 13.1 170.0 167.0
Notebook 3 2.1 158.0 155.5
Other Languages 5 3.8 213.2 210.0
Total 28 19.0 194.5 189.9
Appendix D RM Finetuning Dataset
D.1 Mathematical Reasoning
The RM finetuning for mathematical reasoning uses the MathShepherd dataset( (Wang et al., 2023b)), which contains 444k query-response samples, with some queries having multiple distinct responses. We divide the dataset into a 400k training set and a 44k test set. For RM finetuning, we construct preference pairs by selecting both correct and incorrect responses for the same query. To form the 4.3k test set, we combine one positive and negative sample for each query from the original test set.
We also create two training sets of different sizes: MathShepherd-preference-800k and MathShepherd-preference-40k. The 800k training set is built by combining multiple positive and negative samples for each query in the original training set, resulting in 800k samples. In contrast, the 40k training set randomly selects one positive-negative pair for each query, totaling 40k samples.
D.2 Logical Reasoning
D.2.1 Reclor
Reclor is a human-annotated reading comprehension reasoning dataset, where each sample consists of a passage, a question, and multiple options. To create preference pairs, we combine the correct and incorrect options for the same question. This process results in a total of 14.5k preference pairs, with 14k pairs used for training and 1.5k for testing, forming the Reclor-preference dataset.
D.2.2 LogiQA2.0
For logical reasoning finetuning scaling analysis, we synthesize a preference dataset based on the logical reasoning dataset LogiQA2.0. LogiQA2.0 is a reading comprehension benchmark requiring discrete reasoning over passages, with 96k crowdsourced, adversarially created questions. To answer correctly, models must resolve references (which may point to multiple locations in the input) and perform discrete operations like addition, counting, or sorting. We use four models (Qwen2-7B-Instruct [^1] https://huggingface.co/Qwen/Qwen2-7B-Instruct, Qwen2-72B-Instruct [^2] https://huggingface.co/Qwen/Qwen2-72B-Instruct, DeepSeek-V2-Chat [^3] https://huggingface.co/deepseek-ai/DeepSeek-V2-Chat, and DeepSeek-Coder-V2-Instruct [^4] https://huggingface.co/deepseek-ai/DeepSeek-Coder-V2-Instruct) to sample queries from the LogiQA2.0 dataset multiple times, with sampling $topp=1$, and $topk\in\{0,0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1.0\}$. Answer correctness is annotated using DeepSeek-7B-Math-Compare-Answer [^5] https://huggingface.co/Tianqiao/DeepSeek-7B-Math-Compare-Answer. Correct and incorrect answers are combined to create preference pairs, resulting in 1,019k pairs, with 977k used for training and 42k for testing, forming the LogiQA2.0 preference dataset for RM scaling analysis.
D.3 CodeUltraFeedback_binarized
CodeUltraFeedback_binarized [^6] https://huggingface.co/datasets/coseal/CodeUltraFeedback_binarized is a preference dataset in the code domain, consisting of 9.5k preference pairs. We randomly split the dataset, using 90% of the samples for finetuning training and 10% for testing RM accuracy.
Appendix E More Experimental Results
Due to the length of the paper, we only report important experimental results in the experiment and ablation sections. In this section, we present complementary results in more detail.
E.1 RM Sample Efficiency Comparison: RM accuracy results
(a) MathShepherd-pair / 1.5B
(b) Reclor-pair / 1.5B
(c) LogiQA2.0-pair / 1.5B
(d) MathShepherd-pair / 7B
(e) Reclor-pair / 7B
(f) LogiQA2.0-pair / 7B
Figure 10: Comparison of sample efficiency of RM finetuning: Trends of RM accuracy with sample size increases. Note: The horizontal axis increases exponentially with $\sqrt{2}$. Therefore, the further the data intervals, the more RM data is effectively saved. We use different colors to highlight the results of different model sizes.
We use RM accuracy evaluation to compare the sample efficiency of RM finetuning, as shown in Figure 10. Consistent with the conclusion from the main experiment, CodePMP improves sample efficiency. Under the same sample conditions, it consistently delivers stable improvements, making RM training more efficient.
E.2 Ablation Results on logical Reasoning
Table 5 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for different Model-generated pair construction methods for CodePMP. Overall, the $6.7B\&1.3B$ setting is the best among them.
Table 6 shows the comparison of the BoN accuracy performance of ReClor and LogiQA2.0 using GitHub source code and web crawled code data to build CodePMP training pairs. The results verify that Github code is a better source for pair construction.
Table 7 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for CodePMP models with and without EOC. In general, the setup without using EOC brings better results.
Table 8 shows the comparison of ReClor and LogiQA2.0 BoN accuracies for WCD and WSD lr scheduler settings. The results below show that WSD lr scheduler brings better results than WCD.
Table 5: Comparison of ReClor and LogiQA2.0 BoN accuracies for different Model-generated pair construction methods for CodePMP. Overall, the $6.7B\&1.3B$ setting is the best among them.
Test Set
RM Finetune Constructions Method ReClor LogiQA2.0
ReClor-pair Source Code & 1.3B 0.550 0.4464
Source Code & 1.3B-Des-Clip 0.600 0.4522
Source Code & 6.7B 0.578 0.4605
1.3B & 1.3B-Des-Clip 0.572 0.4745
6.7B & 1.3B-Des-Clip 0.564 0.4809
6.7B & 1.3B 0.608 0.5032
LogiQA2.0-pair Source Code & 1.3B 0.708 0.5217
Source Code & 1.3B-Des-Clip 0.704 0.5268
Source Code & 6.7B 0.714 0.5198
1.3B & 1.3B-Des-Clip 0.738 0.5402
6.7B & 1.3B-Des-Clip 0.742 0.5274
6.7B & 1.3B 0.734 0.5415
Table 6: Comparison of the BoN accuracy performance of ReClor and LogiQA2.0 using GitHub source code and web crawled code data to build CodePMP training pairs. The results below show that Github code is a better source for pair construction.
Test Set
RM Finetune Data Source ReClor LogiQA2.0
ReClor-pair Webpage 0.574 0.4898
Github 0.582 0.4981
LogiQA2.0-pair Webpage 0.742 0.5293
Github 0.752 0.5504
Table 7: Comparison of ReClor and LogiQA2.0 BoN accuracies for CodePMP models with and without EOC. In general, the setup without using EOC brings better results.
Test Set
RM Finetune Model ReClor LogiQA2.0
ReClor-pair w/o EOC 0.582 0.4981
w/ EOC 0.596 0.4617
LogiQA2.0-pair w/o EOC 0.752 0.5504
w/ EOC 0.686 0.4809
Table 8: Comparison of ReClor and LogiQA2.0 BoN accuracies for WCD and WSD lr scheduler settings. The results below show that WSD lr scheduler brings better results than WCD.
Test Set
RM Finetune Model ReClor LogiQA2.0
ReClor-pair WCD 0.552 0.4828
WSD 0.582 0.4981
LogiQA2.0-pair WCD 0.748 0.5204
WSD 0.752 0.5504
Table 9: Cross-validation evaluation: ReClor and LogiQA2.0 BoN accuracies, black numbers are the results of cross-validation. In the cross-validation evaluation, CodePMP can still bring stable improvements on different test sets.
Test Set
RM Finetune Model PMP ReClor LogiQA2.0
ReClor-pair 1.5B 0.534 0.4592
0.582 0.4981
7B 0.724 0.5453
0.756 0.5835
LogiQA2.0-pair 1.5B 0.710 0.5293
0.752 0.5504
7B 0.748 0.6371
0.794 0.6779
E.3 Cross-validations on logical Reasoning
We further conduct cross-dataset validation on logical reasoning. Specifically, we evaluate the RM models finetuned on the ReClor-pairs dataset using the LogiQA2.0 test set. Similarly, we evaluate the RM models finetuned on the LogiQA2.0-pairs dataset using the ReClor test set. As shown in Table 9, CodePMP consistently improves RM evaluation performance, demonstrating that the enhancements CodePMP brings to RM training are robust and generalizable. Note that the BoN accuracies of the RM trained with LogiQA2.0-pair on the Reclor test set are higher than those of the RM trained directly on the Reclor-pair, because the LogiQA2.0-pair dataset is three times larger than the Reclor-pair dataset.
@@ Appendix F Logical Reasoning Evaluation Examples
Table 10: Examples from the Reclor test set, which consists of multiple-choice questions based on a given passage.
ID Text Options Answer
12824 Mayor: Four years ago, when we reorganized the city police department in order to save money, critics claimed that the reorganization would make the police less responsive to citizens and would thus lead to more crime. The police have compiled theft statistics from the years following the reorganization that show that the critics were wrong. There was an overall decrease in reports of thefts of all kinds, including small thefts.
Question: Which of the following, if true, most seriously challenges the mayor’s argument?
1. In other cities where police departments have been similarly reorganized, the numbers of reported thefts have generally risen following reorganization. 2. When city police are perceived as unresponsive, victims of theft are less likely to report thefts to the police. 3. The mayor’s critics generally agree that police statistics concerning crime reports provide the most reliable available data on crime rates. 4. The mayor’s reorganization of the police department failed to save as much money as it was intended to save. 1
218 Jupiter is a gas giant planet and the largest planet in the solar system. Its mass is 2.5 times the total mass of the other seven planets in the solar system. Observations have found that most of the more than 70 moons surrounding Jupiter are composed of water ice. Therefore, Jupiter’s atmosphere should contain a considerable amount of water.
Question: Which of the followings, if true, can best support the above statement?
1. After hundreds of millions of years, the satellite may slowly fall onto the planet. 2. Many of the water in interstellar space exists in gaseous form. 3. Uranus is also a gas giant planet, and it has been confirmed that it contains a lot of water ice. 4. The satellite and the planets around it were formed from the same gas and dust at the same time. 3
10376 Lake Dali is a barrier lake on the plateau formed by volcanic eruptions. Like salmon living in the sea, Hua Zi fish-Leuciscus waleckii, which lives in a brackish lake, must migrate to the upper reaches of the Tanshui River to spawn and breed, although the four rivers currently flowing into Lake Dali are inland rivers, and none of them leads to the sea. Scientists are still convinced that the Huaziyu in Lake Dali first migrated from the ocean.
Question: Which of the following options, if true, provides the best explanation for scientists’ beliefs?
1. The Leuciscus waleckii that lives in the waters such as Heilongjiang is twice as big as the Leuciscus waleckii fish in Lake Dari. 2. The caught Hua Zi fish can only survive for a day or two after being put into sea water or fresh water, and will decay quickly after death. 3. Melting glaciers will form Lake Dali, and the overflowing lake was once connected to the Liao River, which flowed into the ocean. 4. The researchers put the fry of Hua Zi fish in Dali Lake into Gainao thousands of miles away, and the culture was successful. 2
13334 It is repeatedly claimed that the dumping of nuclear waste poses no threat to people living nearby. If this claim could be made with certainty, there would be no reason for not locating sites in areas of dense population. But the policy of dumping nuclear waste only in the more sparsely populated regions indicates, at the very least, some misgiving about safety on the part of those responsible for policy.
Question: Which one of the following, if true, would most seriously weaken the argument?
1. Until there is no shred of doubt that nuclear dumps are safe, it makes sense to situate them where they pose the least threat to the public. 2. There are dangers associated with chemical waste, and it, too, is dumped away from areas of dense population. 3. In the event of an accident, it is certain that fewer people would be harmed in a sparsely populated than in a densely populated area. 4. Dumping of nuclear waste poses fewer economic and bureaucratic problems in sparsely populated than in densely populated areas. 3

### v2

@@ Appendix A Hyperparameters
We outline key hyperparameters in this section. In the tables, WSD refers to the warmup-stable-decay learning rate scheduler (Hu et al., 2024), which has the benefit of reducing the time required for scaling law experiments. Specifically, Table 7 lists the hyperparameters for CodePMP training, Table 8 details those for mathematical reasoning RM fine-tuning, Table 9 covers logical reasoning RM fine-tuning, and Table 10 presents the hyperparameters for BON generation.
Hyperparameter Qwen2-1.5B Qwen2-7B
epoch 1 1
batch size 1024 1024
learning rate 3e-6 1e-6
lr scheduler WSD WSD
warmup ratio 0.03 0.03
decay ratio 0.1 0.1
weight decay 0.1 0.1
max length 1024 1024
Table 7: CodePMP training hyperparameters.
Hyperparameter Qwen2-1.5B Qwen2-7B
epoch 1 1
batch size 64 64
learning rate 1e-6 3e-7
lr scheduler WCD WCD
warmup ratio 0.03 0.03
weight decay 0 0
max length 1024 1024
Table 8: Mathematical reasoning RM finetuning hyperparameters.
Hyperparameter Qwen2-1.5B Qwen2-7B
epoch 1 1
batch size 64 64
learning rate 1e-5 1e-5
lr scheduler WCD WCD
warmup ratio 0.25 0.25
weight decay 0 0
max length 1024 1024
Table 9: Logical reasoning RM finetuning hyperparameters.
Hyperparameter MetaMath-Mistral-7B
temperature 0.7
top-p 1
Table 10: Best-of-N generation hyperparameters.
Appendix B CodePMP Dataset Statistics
Table 11 presents the average token lengths of responses in the CodePMP dataset. The similar lengths between chosen and rejected responses (194.5 vs. 189.9 tokens) ensure that response length does not introduce bias in the learning process. The dataset comprises 28 million files totaling 19 billion tokens, with Python (13.1B tokens), Jupyter Notebooks (2.1B tokens), and other languages (3.8B tokens) providing diverse coverage that facilitates model generalization.
Language Chosen Rejected
Python 170.0 167.0
Notebook 158.0 155.5
Other Languages 213.2 210.0
Total 194.5 189.9
Table 11: Average token lengths of responses in the CodePMP dataset by language category.
(a) MATH-shepherd / 1.5B
(b) MATH-shepherd / 7B
(c) Reclor+Logiqa / 1.5B
(d) Reclor+Logiqa / 7B
Figure 8: Comparison of sample efficiency of RM fine-tuning: Trends of RM accuracy with sample size increases.
(a) MATH-shepherd / 1.5B
(b) MATH-shepherd / 7B
(c) Reclor+Logiqa / 1.5B
(d) Reclor+Logiqa / 7B
Figure 9: Comparison of sample efficiency of RM fine-tuning: Trends of Multi-choice accuracy or Best-of-4 with sample size increases.
Appendix C Further Comparisons and Cross-Domain Evaluations
C.1 Comparison with Majority Voting
We compare CodePMP with a majority-voting baseline under the same experimental setup on GSM8K and MATH. Table 12 shows that CodePMP outperforms majority voting, especially on more complex tasks like MATH.
Method GSM8K MATH
CodePMP 0.8484 0.41
Majority Voting 0.8453 0.37
Table 12: Comparison of CodePMP and majority voting on GSM8K and MATH.
C.2 Sample Efficiency Improvements on Reclor and LogiQA
We finetune the RM on preference pairs using only Reclor or LogiQA and then evaluate them on their respective test sets. As shown in Figures 8 and 9, PMP demonstrates a clear advantage in sample efficiency, reflected in both RM accuracy and Best-of-N evaluation. The results reveal that even with substantially fewer training samples, reward models initialized with CodePMP achieve comparable or better performance than models trained from scratch with many more samples, highlighting the significant sample efficiency benefits of our approach for logical reasoning tasks.
C.3 Performance on Coding Tasks
We evaluate CodePMP’s effectiveness on actual code generation tasks by conducting two types of evaluations: reward model accuracy assessment and code generation evaluation.
First, we assess the reward model’s accuracy on the CodeUltraFeedback benchmark, which consists of preference pairs in the code domain. We fine-tuned Qwen2 models on the CodeUltraFeedback_binarized dataset (8.5k preference pairs), both with and without CodePMP initialization. Table 13 presents the accuracy results across model sizes.
Model PMP CodeUltraFeedback
Accuracy
1.5B 0.6841
0.758
7B 0.6912
0.7619
Table 13: Performance on CodeUltraFeedback benchmark shows that CodePMP improves in-domain code reward modeling.
As shown in Table 13, reward models initialized with CodePMP consistently outperform those without PMP initialization on the CodeUltraFeedback benchmark. For the 1.5B model, CodePMP initialization improves accuracy from 0.6841 to 0.758, while for the 7B model, accuracy increases from 0.6912 to 0.7619. These results demonstrate that CodePMP effectively enhances reward models’ ability to evaluate code quality.
Beyond reward model evaluation, we also assess whether this improved evaluation capability translates to better code generation outcomes using the HumanEval benchmark. For this evaluation, we used deepseek-coder-6.7b-instruct as the generator and Qwen2-7B as the reward model (RM). We fine-tuned the RM on the same CodeUltraFeedback_binarized dataset, both with and without CodePMP initialization. Table 14 presents Pass@1 (0-shot) results under different $N$ values.
BoN Qwen2-7B w/o PMP Qwen2-7B w/ PMP N=1 0.7134 0.7134 N=2 0.7317 0.7195 N=4 0.7073 0.7622 N=8 0.6890 0.7683 N=16 0.6951 0.7256 N=32 0.6585 0.7378 N=64 0.6829 0.7134 N=128 0.6707 0.7012 N=256 0.6707 0.7195
Table 14: HumanEval results (Pass@1, 0-shot) for different numbers of sampled solutions $N$. The generator is deepseek-coder-6.7b-instruct.
The results in Table 14 indicate that CodePMP initialization provides a generally more stable and higher-accuracy selection mechanism compared to direct training, especially as $N$ varies. For most values of $N$, the model with CodePMP initialization achieves better Pass@1 scores, with particularly notable improvements at $N=4$ (0.7622 vs. 0.7073), $N=8$ (0.7683 vs. 0.6890), and $N=32$ (0.7378 vs. 0.6585). Without CodePMP, we observe performance degradation at higher $N$ values, while CodePMP-initialized models maintain more consistent performance. This finding is particularly significant since HumanEval evaluates actual code generation rather than just preference prediction, demonstrating that the benefits of CodePMP extend beyond improved preference modeling to better code generation outcomes.
Summary
These additional experiments demonstrate that CodePMP:
  • Outperforms majority voting in both simpler (GSM8K) and more challenging (MATH) settings.
  • Demonstrates significant sample efficiency improvements on logical reasoning tasks (Reclor and LogiQA), with models initialized with CodePMP achieving better performance with fewer training samples.
  • Provides more stable and accurate code evaluation on HumanEval, showing benefits for practical code generation tasks.
Thus, CodePMP provides a scalable and effective approach to improving large language models across different domains and tasks.
Appendix D Comprehensive Data Diversity Analysis
To validate the quality of our synthetic data, we conducted comprehensive diversity analyses using established methodologies from the research on synthetic text data generation (Zhu et al., 2024). These analyses aim to demonstrate that our synthetic data maintains sufficient diversity while effectively capturing the distributions present in human-generated data.
D.1 N-gram Feature Distribution Analysis
We mapped text n-gram features to fixed hash buckets (100 buckets) and analyzed their distribution patterns to measure lexical diversity. Figures 10 and 11 show the comparison between human-generated data and our synthetic data.
Figure 10: Unigram distribution comparison (left: distribution for human data, right: distribution for synthetic data).
Figure 11: Bigram distribution comparison (left: distribution for human data, right: distribution for synthetic data).
Table 15 presents the density values for unigram and bigram distributions across different data sources.
Data Source Unigram Bigram Human 134,538.40 133,538.41 Strong Model 97,653.69 96,653.70 Weak Model 93,691.69 92,691.70
Table 15: N-gram density values for human and synthetic data.
The distribution graphs show that, compared to human data, our synthetic data has more uniform n-gram distributions, without the concentration peaks common in synthetic data. The density values further quantify this advantage—synthetic text’s n-gram density values (Strong Model: 97,653.69, Weak Model: 93,691.69) are significantly lower than human text (134,538.40), demonstrating more balanced distribution across hash buckets.
D.2 Embedding Space Visualization
To further evaluate the semantic diversity of our synthetic data, we mapped semantic features of both human and synthetic data to a 2D space, as shown in Figure 12.
Figure 12: Embedding space visualization (left: distribution for human data, right: distribution for synthetic data).
Both synthetic and human data show wide and dispersed distributions in the embedding space with highly overlapping distribution ranges, indicating our synthetic data captures a similarly broad semantic space as human data.
D.3 KL Divergence Analysis
We quantified the distribution differences between synthetic and human data using KL divergence to evaluate how closely our synthetic data approximates natural distributions. Table 16 presents these results.
N-gram Human Internal Strong Model Weak Model (Bootstrap) vs Human vs Human 1-gram 0.2502 0.4290 0.4631 2-gram 0.6904 1.3500 1.4281 3-gram 1.3012 2.5660 2.6693
Table 16: KL divergence values comparing different data distributions.
These results demonstrate that the distribution differences between our synthetic data and human data fall within acceptable ranges relative to internal human data variation.
D.4 Comprehensive Validation of Synthetic Preference Data
Our synthetic data generation approach relies on two key assumptions: (1) larger models from the same family produce higher-quality code than smaller ones, and (2) this quality difference creates consistent preference signals suitable for training. We conducted both theoretical and empirical validation to confirm these assumptions.
D.4.1 Validation of Strong-Weak Model Ability Differences
To validate our first assumption, we analyzed ability differences between strong and weak models across various code-related dimensions.
(a) Same model family
(b) Different model families
Figure 13: Radar charts showing ability differences between strong and weak LLMs across various dimensions.
Figure 13(a) provides strong evidence supporting our assumption: when using models from the same architectural family with different parameter counts, the stronger model consistently outperforms the weaker model across all ability dimensions. This uniform superiority ensures that synthetic preference pairs have clear and consistent quality differences, creating reliable signals for training preference models.
For comparison, Figure 13(b) shows what happens when models from different families are paired. Here, we observe irregular and inconsistent differences, with some dimensions showing negligible gaps or even inversions. Such inconsistencies could potentially introduce noise into the preference signals, undermining training data quality.
D.4.2 External Evaluation of Preference Consistency
Having confirmed the underlying ability differences, we next validated whether these differences translate to consistent preference judgments. We conducted preference annotation experiments using GPT-4o to evaluate the consistency of preferences in our dataset.
The results show that our synthetic CodePMP data achieved a preference consistency rate of 75.12%. This is notably higher than the more costly CodeUltraFeedback preference dataset (71.56%), demonstrating that the preference distinction in our synthetic data (based on our assumption that "larger models generate better code than smaller models") is sufficiently clear and consistent.
This external validation reinforces our first finding - not only do larger models from the same family consistently outperform smaller ones across all dimensions, but this performance gap is readily detectable by strong evaluator models, resulting in consistent preference judgments.
As CodePMP is fundamentally a pretraining process, we deliberately simplified our assumptions to enable scalable preference data creation with minimal additional validation. Our multi-faceted validation approach confirms that this simple yet effective methodology produces high-quality, consistent preference data suitable for large-scale pretraining.
D.5 Source Data Quality and Experimental Validation
Our method achieves excellent diversity due to the high quality of our source data:
  • We collected over 130 million code snippets from GitHub, covering all common programming languages and task types on open-source platforms, ensuring breadth and depth in our source data.
Furthermore, our experimental results validate the effectiveness of our synthetic data diversity:
  • As shown in Figure 6(a), our synthesis strategy outperforms preference pairs constructed directly from source code, indirectly proving that our synthesis process enhances data diversity and quality.
This comprehensive diversity analysis confirms that our synthetic data generation approach produces high-quality, diverse data that effectively captures the distribution characteristics of human-generated code. The balanced distribution patterns and semantic coverage demonstrate that our synthetic data is well-suited for training robust reward models.
Appendix E Detailed Implementation of CodePMP
In this section, we provide a detailed overview of the Code Preference Model Pretraining (CodePMP) implementation. The following description illustrates the systematic process of generating and utilizing code-preference pairs for pretraining preference models, which can then be fine-tuned for downstream reasoning tasks.
The algorithm begins with a source code repository, a strong CodeLLM (in our implementation, deepseek-coder-6.7b-instruct), and a weaker CodeLLM (deepseek-coder-1.3b-instruct). First, descriptions are generated for each code snippet in the repository using the strong model. For each description, the strong model generates a high-quality chosen response, while the weaker model generates a less optimal rejected response. These pairs are used to calculate both language modeling loss (on the responses) and reward modeling loss (comparing chosen vs. rejected responses). The final training objective combines these two loss components.
This scalable approach allows for creating millions of preference pairs without expensive human annotation, providing an effective initialization for reward models that will later be fine-tuned on specific reasoning tasks.
Algorithm 1 Code Preference Model Pretraining
0:  Source code repository $S$, Strong CodeLLM $M_{\text{strong}}$, Weak CodeLLM $M_{\text{weak}}$
0:  Pretrained Model
  Input: Source code $S$
  Summarize description $D$ using $M_{\text{strong}}$ on $S$
  for each $D_{i}\in D$ do
     Generate Chosen Response using $M_{\text{strong}}$
     Generate Rejected Response using $M_{\text{weak}}$
  end for
  Calculate LM Loss $\mathcal{L}_{\text{LM}}$ on Response
  Calculate RM Loss $\mathcal{L}_{\text{RM}}$ using Chosen Response and Rejected Response
  Train PMP Model using $\mathcal{L}_{\text{PMP}}=\mathcal{L}_{\text{RM}}+\mathcal{L}_{\text{LM}}$
@@ Appendix F Logical Reasoning Evaluation Examples
Table 17: Examples from the Reclor test set, illustrating multiple-choice questions format with passages and questions.
ID Passage Question Ans.
12824 Mayor: When we reorganized the police department, critics claimed it would make police less responsive and lead to more crime. Statistics show an overall decrease in thefts after reorganization. Which statement most challenges the mayor’s argument? (1) Similar reorganizations in other cities led to increased thefts. (2) Unresponsive police reduce theft reporting rates. (3) Critics agree police statistics are reliable. (4) The reorganization saved less money than planned. 2
218 Jupiter is the largest planet with mass 2.5 times that of all other planets combined. Most of Jupiter’s 70+ moons are water ice. What best supports that Jupiter’s atmosphere should contain water? (1) Satellites may eventually fall onto planets. (2) Interstellar water exists as gas. (3) Uranus, also a gas giant, contains water ice. (4) Satellites and planets form from the same materials. 3
10376 Lake Dali fish must migrate to river headwaters to breed, though no rivers connect to the sea. Scientists believe these fish originally came from the ocean. What best explains scientists’ belief? (1) Similar fish elsewhere are larger. (2) The fish quickly die in sea/fresh water. (3) Lake Dali was once connected to an ocean-bound river. (4) Fish from Lake Dali survived in far-away lakes. 2
13334 If nuclear waste posed no threat, it could be placed in populated areas. But it is only dumped in sparsely populated regions, suggesting safety concerns. What would most weaken this argument? (1) Uncertain safety justifies minimal risk placement. (2) Chemical waste is also dumped away from population. (3) Accidents affect fewer people in sparsely populated areas. (4) Remote locations reduce bureaucratic complications. 3